OpenStudy (anonymous):

A car travels around a horizontal circular track at a constant speed of 58 km/h. If the radius of the track is 43 metres, what is the magnitude of the cars acceleration in m/s2?

OpenStudy (anonymous):

a=v^2 /r plug the values??

OpenStudy (anonymous):

the formula is a=v^2/r so by subtituting the values a=(58)^2/43 a= 78.23m/s^2

OpenStudy (anonymous):

the answer which the book gives is 6.0m/s^2 and i dont know how to get this value.

OpenStudy (anonymous):

@sgholap you forgot to convert velocity to m/s

OpenStudy (anonymous):

@michael convert velocity into m/s and the plug the values!!! you are sure to get it??

OpenStudy (anonymous):

ohhh!!!! i am really sorry michael i forget to convert it just do wat sarkar said will get the answer right